### v2

@@ 3 Quantum Computing Lab, Academic Computer Center Cyfronet AGH, Nawojki 11, Krakow, 30-950, Poland
Samplers: Through Samplers we specify the method by which we want to detect communities, providing the necessary parameters. Currently, several classical solutions are implemented, such as Gurobi, Louvain [10], Bayan[15], Leiden [12], hybrid – using DQM [25], and purely quantum – using Advantage [19, 65]. We divided samplers based on how the divisions are done;
@@
Although Samplers provide community detection methods and can perform the community divisions themselves, the Seacher and IterativeSearcher are the dedicated tools for the user to perform the community detection process comprehensively.
@@
High-level Python code examples to use the quantum resources for modularity maximization and interpretation

### v3

@@ Acknowledgments
Samplers: Through Samplers we specify the method by which we want to detect communities, providing the necessary parameters. Currently, several classical solutions are implemented, such as Gurobi, Louvain [Blondel2008], Bayan[aref2023], Leiden [Traag2019], Infomap [Rosvall2008], SBM Inference [Peixoto2014], Spectral clustering, hybrid – using DQM [Wierzbinski2023], and purely quantum – using Advantage [Johnson2011, lamza_qhyper_2024]. We divided the samplers based on how the divisions are done;
@@
Although Samplers provide community detection methods and can perform community divisions themselves, the Seacher and IterativeSearcher are dedicated tools that enable the user to perform the community detection process comprehensively.
@@
S7.2 High-level Python code examples to use the quantum resources for modularity maximization and interpretation